A race car accelerates uniformly from 18.5 m/s to 46.1 m/s in 2.47 seconds. Determine the acceleration of the car and the distance traveled.
1
Expert's answer
2020-05-21T13:13:40-0400

When the car accelerates uniformly, the speed increases linearly: "v(t) = v_0 + a t", from where given "v_0 = 18.5 \frac{m}{s}" and "v(2.47) = 46.1\frac{m}{s}" the acceleration is "a = \frac{v(2.47) - v_0}{2.47} \approx 11.17\frac{m}{s^2}".

The law of motion with constant acceleration is "s(t) = v_0 t + \frac{a t^2}{2}", and substituting "a = \frac{v(t) - v_0}{t}", obtain "s(t) = v_0 t + \frac{v(t) - v_0}{2} t \approx 79.8 m".